OpenStudy (anonymous):

find (z-8)^4/3=2 Where z is a real number...

OpenStudy (anonymous):

\[(z-8)^{4/3}=2 \rightarrow (z-8)=2^{3/4} \rightarrow z-8=\sqrt[4]{2^{3}}\] i.e. \[z-8=\sqrt[4]{8} \rightarrow z=8+\sqrt[4]{8}\]